# Architectural Co-Design for Zero-Shot Anomaly Detection: Decoupling Representation and Dynamically Fusing Features in CLIP

### v1

@@ Abstract
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: Vision and text encoders are partitioned into hierarchical Groups. Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@ 1 Introduction
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are:
  • A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation.
  • A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy.
  • Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
@@ 2.1 Hierarchical Feature Adaptation with Local Priors
The adapter features a parallel multi-branch design ($3\times 3$ and $5\times 5$ convolutions) operating within a LoRA bottleneck. This multi-scale convolutional design allows the adapter to capture local patterns at varied receptive fields, which is crucial for identifying anomalies of different sizes. For each branch $k$, the input $X_{\text{in}}$ is projected to a low-rank tensor $X_{\text{proj}}$, reshaped into a spatial format ($\mathcal{R}_{\text{2D}}$), and passed through a convolutional bottleneck comprising a strided convolution ($\text{Conv}_{\text{down}}^{(k)}$) and a transposed convolution ($\text{Conv}_{\text{up}}^{(k)}$).
$X_{\text{proj}}=X_{\text{in}}W_{\text{down}}$ (1)
$X_{\text{up}}^{(k)}=\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}(\operatorname{Conv}_{\text{down}}^{(k)}(\mathcal{R}_{\text{2D}}(X_{\text{proj}})))$ (2)
The output $X_{\text{up}}^{(k)}$ is reshaped back to sequence format ($\mathcal{R}_{\text{seq}}$) and projected by $W_{\text{up}}$. The outputs from all branches are then concatenated and fused by a $1\times 1$ convolution to generate the residual update $\Delta X$, which enhances the main vision stream.
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}(X_{\text{up}}^{(k)})W_{\text{up}}$ (3)
$X_{\text{cat}}=\operatorname{Concat}(X_{\text{branch\_out}}^{(3)},X_{\text{branch\_out}}^{(5)})$ (4)
$\Delta X=\mathcal{R}_{\text{seq}}(\operatorname{Conv}_{1\times 1}(\mathcal{R}_{\text{2D}}(X_{\text{cat}})))$ (5)
@@ 2.2 Dynamic Fusion Gateway
To overcome static fusion, our Dynamic Fusion Gateway (DFG) employs a vision-guided mechanism to generate custom semantic descriptors $(T_{i}^{N},T_{i}^{A})$ for each vision feature group $V_{i}$. A global context vector is extracted from $V_{i}$ and passed through an $\text{MLP}_{\text{gate}}$ to produce logits over all $N$ text feature levels. This context vector acts as a query to dynamically weigh the importance of all text levels, effectively creating a vision-guided attention mechanism over the text features. These logits are normalized via Softmax to generate dynamic weight vectors, $\bm{\omega}_{i}^{N}$ and $\bm{\omega}_{i}^{A}$:
$\bm{\omega}_{i}^{N}=[\omega_{i,1}^{N},\dots,\omega_{i,N}^{N}]=\text{Softmax}(L_{i}^{N})$ (6)
$\bm{\omega}_{i}^{A}=[\omega_{i,1}^{A},\dots,\omega_{i,N}^{A}]=\text{Softmax}(L_{i}^{A})$ (7)
These weights dynamically fuse the set of multi-level text features $\{(T_{j}^{N},T_{j}^{A})\}_{j=1}^{N}$ to produce context-aware ”normal” ($T_{i}^{N}$) and ”abnormal” ($T_{i}^{A}$) descriptors:
$T_{i}^{N}=\sum_{j=1}^{N}\omega_{i,j}^{N}\cdot T_{j}^{N}$ (8)
$T_{i}^{A}=\sum_{j=1}^{N}\omega_{i,j}^{A}\cdot T_{j}^{A}$ (9)
The final anomaly map is the average of level-specific maps $\{M_{i}\}$, each computed from the cosine similarity between $V_{i}$ and its corresponding descriptor pair.
@@ 2.3 Training Objective and Inference
The framework is trained end-to-end with a composite loss function.
Segmentation Loss. Level-specific anomaly maps $\{M_{i}\}$ are averaged to form the final prediction map $M_{\text{seg}}$, which is optimized using a combined Focal and Dice loss $\mathcal{L}_{\text{seg}}$ against the ground-truth mask $M_{\text{GT}}$:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{F}}L_{\text{F}}(M_{\text{seg}},M_{\text{GT}})+\lambda_{\text{D}}L_{\text{D}}(M_{\text{seg}},M_{\text{GT}})$ (10)
Classification Loss. For image-level classification, we compute the cosine similarity between the final global visual feature ($V_{N}^{\text{cls}}$) and the original, unfused text features from the last group ($T_{N}^{N},T_{N}^{A}$), supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$. We use the original text features here to provide a stable, high-level semantic anchor for the classification task, decoupling it from the fine-grained fusion used for segmentation.
Total Objective. The final training objective is a weighted sum of the two losses:
@@
where $\lambda_{\text{F}},\lambda_{\text{D}},\text{ and }\lambda_{\text{cls}}$ are balancing hyperparameters.
@@ 3 Experiments
We conduct extensive experiments to validate the effectiveness of our proposed Architectural Co-Design with Dynamic Fusion (ACD-CLIP). We benchmark our method against state-of-the-art (SOTA) approaches on multiple industrial and medical datasets and perform detailed ablation studies to analyze the contribution of each component.
@@ 3.1 Experimental Setup
Datasets and Baselines. We evaluate on a wide range of public benchmarks. For the industrial domain, we use MVTec-AD [17], BTAD [18], MPDD [19], RSDD [20], and VisA [21]. For the medical domain, we use BrainMRI, Liver CT, and Retina OCT from BMAD [22], ColonDB [23], ClinicDB [24], Kvasir [25] to demonstrate cross-domain generalization. We compare our method against several recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8]. To maintain a strict zero-shot setting, our model is trained on the VisA dataset for all benchmarks except VisA itself, as there is no category overlap. For evaluation on the VisA benchmark, the model is trained exclusively on the MVTec-AD dataset.
Metrics. We follow standard evaluation protocols, reporting the Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both Pixel-level segmentation and Image-level classification tasks.
3.2 Main Results
Our method demonstrates superior performance across diverse ZSAD benchmarks. As shown in Table 1, on the MVTec-AD benchmark, our model achieves a pixel-level AUROC/AP of 91.7%/44.0% and an image-level AUROC of 92.4%, significantly outperforming prior methods like AnomalyCLIP (91.1%/34.5% pixel; 90.3% image). This advantage is consistent across other industrial datasets such as BTAD and VisA. Crucially, our approach demonstrates strong generalization to the medical domain. For instance, on ClinicDB, our model obtains a 90.4% pixel-level AUROC, surpassing AnomalyCLIP (85.9%) by a large margin. Qualitative results (Figure 3) corroborate these findings, showing our model generates precise anomaly maps that align closely with the ground truth, exhibiting higher fidelity and less background noise than competing methods.
@@ 3.2 Main Results
Table 2: Ablation study of our proposed components. We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
@@
3.3 Ablation and Analysis
To isolate the effects of our components, we conduct an ablation study on the MVTEC-AD dataset with a two-group configuration ($N=2$). We validate our two core contributions, the Conv-LoRA Adapter and the Dynamic Fusion Gateway (DFG), against a baseline that uses a standard linear adapter and static fusion. As shown in Table 2, both components individually yield significant improvements. Specifically, adding the Conv-LoRA Adapter boosts the pixel-level AUROC by 6.8% (to 89.1%), confirming the benefit of injecting local inductive biases for dense prediction. Concurrently, the DFG enhances the image-level AUROC by 4.7% (to 85.9%), demonstrating the effectiveness of its vision-guided, context-aware fusion. Our full model, which integrates both components, achieves the highest performance (91.7% pixel-level and 90.9% image-level AUROC), proving a strong synergistic effect between the two modules. These results validate that both components are essential and complementary for achieving a substantial performance uplift.

### v2

@@ Abstract
Fig. 2: Overview of the proposed ACD-CLIP architecture. (a) The Overall Framework: We structure CLIP’s vision and text encoders into a hierarchy NN sequential Groups (as illustrated, $N=3$). Each vision group is enhanced by a trainable Conv-LoRA Adapter to instill local priors. The Dynamic Fusion Gateway (DFG) then uses each visual feature $V_{i}$ to generate a tailored text descriptor for producing a level-specific anomaly map $M_{i}$. (b) The Conv-LoRA Adapter: Our parameter-efficient adapter features a multi-branch design with multi-scale convolutions inside a LoRA bottleneck.
@@ 1 Introduction
To address these core issues, we introduce the Architectural Co-Design CLIP (ACD-CLIP) framework [Figure 1(b)], a novel approach that synergistically refines both feature representation and cross-modal fusion. The synergy is crucial: our adapter provides the fine-grained visual details that static fusion would overlook, while the gateway supplies the adaptive mechanism required to fully exploit them. Our architectural modifications are enabled by Parameter-Efficient Fine-Tuning (PEFT) [13, 14]. While recent works have shown the value of integrating convolutional structures into PEFT adapters for general vision tasks [15, 16], our work is the first to co-design such an adapter with a dynamic fusion mechanism specifically for ZSAD. Our contributions are: 1) A parameter-efficient Conv-LoRA Adapter to inject local inductive biases for fine-grained representation. 2) A Dynamic Fusion Gateway, a vision-guided mechanism enabling a flexible, bidirectional fusion policy. 3) Validation of our synergistic co-design, which yields significant performance gains on diverse ZSAD benchmarks.
@@ 2.1 Hierarchical Feature Adaptation with Local Priors
The adapter features a parallel multi-branch design, where each branch employs a distinct convolutional kernel size $k\times k$ for $k\in\{3,5\}$, operating within a LoRA bottleneck. This allows it to capture local patterns at varied receptive fields. The entire operation of the adapter, denoted as $\Phi_{\text{Adapter}}$, is formulated to generate a residual update $\Delta X$ as follows:
$\Delta X$ $=\Phi_{\text{Adapter}}(X_{\text{in}})$ (1)
$=\mathcal{R}_{\text{seq}}\Big{(}\operatorname{Conv}_{1\times 1}\big{(}\mathcal{R}_{\text{2D}}\big{(}\operatorname{Concat}\big{[}X_{\text{branch\_out}}^{(k)}\big{]}\big{)}\big{)}\Big{)}$
The output of each branch, $X_{\text{branch\_out}}^{(k)}$, is generated in a two-step process. First, the input tensor $X_{\text{in}}$ is passed through the LoRA down-projection and a strided down-convolution to produce an intermediate downsampled feature map, $X_{\text{down}}^{(k)}$:
$X_{\text{down}}^{(k)}=\operatorname{Conv}_{\text{down}}^{(k)}\left(\mathcal{R}_{\text{2D}}(X_{\text{in}}W_{\text{down}})\right)$ (2)
This feature map is then upsampled via a transposed convolution, reshaped, and passed through the LoRA up-projection matrix to yield the final branch output:
$X_{\text{branch\_out}}^{(k)}=\mathcal{R}_{\text{seq}}\left(\frac{1}{k^{2}}\operatorname{Conv}_{\text{up}}^{(k)}\left(X_{\text{down}}^{(k)}\right)\right)W_{\text{up}}$ (3)
Here, $W_{\text{down}}$ and $W_{\text{up}}$ are the LoRA projection matrices. $\mathcal{R}_{\text{2D}}$ and $\mathcal{R}_{\text{seq}}$ are operators for reshaping between sequence and spatial formats. $\operatorname{Conv}_{\text{down}}^{(k)}$ is a strided convolution and $\operatorname{Conv}_{\text{up}}^{(k)}$ is a transposed convolution.
@@ 2.2 Dynamic Fusion Gateway
To overcome static fusion, our Dynamic Fusion Gateway (DFG) generates custom semantic descriptors for each visual feature group $V_{i}$. A global context vector $v_{i}^{\text{global}}=\operatorname{GAP}(V_{i})$ is extracted and passed through a gating MLP to produce logits $\bm{l}_{i}^{s}$ for each semantic state $s\in\{N,A\}$ (Normal, Abnormal). These logits determine the dynamic fusion weights $\omega_{i,j}^{s}$ for the set of all text features $\{T_{j}^{s}\}_{j=1}^{N}$:
$\omega_{i,j}^{s}=\frac{\exp(l_{i,j}^{s})}{\sum_{k=1}^{N}\exp(l_{i,k}^{s})},\quad\text{where }\bm{l}_{i}^{s}=\text{MLP}_{\text{gate}}^{s}(v_{i}^{\text{global}})$ (4)
These weights dynamically fuse the multi-level text features to produce context-aware normal ($T_{i}^{N}$) and abnormal ($T_{i}^{A}$) descriptors:
$T_{i}^{s}=\sum_{j=1}^{N}\omega_{i,j}^{s}\cdot T_{j}^{s},\quad\text{for }s\in\{N,A\}$ (5)
The resulting level-specific anomaly map, $M_{i}$, is then computed by applying a softmax function over the patch-wise cosine similarities between the visual features $V_{i}$ and their corresponding dynamic text descriptors:
$M_{i}=\frac{\exp(\text{sim}(V_{i},T_{i}^{A})/\tau)}{\exp(\text{sim}(V_{i},T_{i}^{N})/\tau)+\exp(\text{sim}(V_{i},T_{i}^{A})/\tau)}$ (6)
where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity operator and $\tau$ is a temperature parameter. The final anomaly map is the average of these level-specific maps $\{M_{i}\}$.
@@ 2.3 Training Objective and Inference
The framework is trained end-to-end with a composite loss function. which jointly addresses pixel-level segmentation and image-level classification.
Segmentation Loss. The final prediction map is the average of level-specific maps, $M_{\text{seg}}=\frac{1}{N}\sum_{i=1}^{N}M_{i}$. It is optimized against the ground-truth mask $M_{\text{GT}}$ using a combined Focal [17] and Dice loss [18]:
$\mathcal{L}_{\text{seg}}=\lambda_{\text{Focal}}\mathcal{L}_{\text{Focal}}(M_{\text{seg}},M_{\text{GT}})$
$+\lambda_{\text{Dice}}\mathcal{L}_{\text{Dice}}(M_{\text{seg}},M_{\text{GT}})$ (7)
where $\lambda_{\text{Focal}}$ and $\lambda_{\text{Dice}}$ are balancing hyperparameters.
Classification Loss. For image-level classification, we compute the cosine similarity between the final global visual feature and the original, unfused text features from the last group, supervised via Cross-Entropy loss $\mathcal{L}_{\text{cls}}$. We use the original text features here to provide a stable, high-level semantic anchor for the classification task.
Total Objective. The final training objective is a weighted sum of the two losses, balanced by the hyperparameter $\lambda_{\text{cls}}$:
@@ 3 Experiments
To demonstrate the efficacy and robustness of our ACD-CLIP framework, we conducted a comprehensive evaluation. This involved benchmarking our model against state-of-the-art (SOTA) competitors across a wide range of datasets and performing in-depth ablation studies to quantify the contribution of each individual component.
@@ 3.1 Experimental Setup
Datasets and Baselines. We perform a comprehensive evaluation on 12 public benchmarks spanning the industrial domain (MVTec-AD [19], VisA [20], BTAD [21], MPDD [22], RSDD [23]) and the medical domain (BrainMRI, Liver CT, and Retina OCT from BMAD [24], ColonDB [25], ClinicDB [26], Kvasir [27]). We benchmark against recent SOTA methods, including WinCLIP [6], CLIP-AD [9], AnomalyCLIP [7], and AdaCLIP [8].
Implementation Details. Our framework is built upon the pre-trained CLIP model with a ViT-L/14 backbone at a 336px resolution. To maintain a strict zero-shot setting, we train our model on an auxiliary dataset with no category overlap (VisA for all non-VisA benchmarks, and MVTec-AD for the VisA benchmark). We report the standard Area Under the Receiver Operating Characteristic curve (AUROC) and Average Precision (AP) for both pixel-level and image-level tasks.
3.2 Main Results and Analysis
As shown in Table 1, our ACD-CLIP framework consistently establishes a new state-of-the-art across both industrial and medical domains. On the widely-used MVTec-AD benchmark, our model demonstrates a clear advantage in fine-grained localization, outperforming methods like AnomalyCLIP in pixel-level AP by a significant margin of nearly +10%. This superiority in capturing precise anomaly boundaries is a direct result of our architectural co-design.
Crucially, our approach shows strong cross-domain generalization. When trained solely on industrial data, ACD-CLIP excels on challenging medical benchmarks, achieving a pixel-level AUROC of 90.4% on ClinicDB and 96.6% on BrainMRI. These quantitative gains are supported by qualitative results (Figure 3), which show our model generates significantly cleaner and more accurate anomaly maps with less background noise compared to other baselines.
@@ 3.2 Main Results and Analysis
Table 2: Ablation study of our proposed components on the MVTec-AD dataset with a two-group configuration ($N=2$). We report the average AUROC (%) for both pixel-level and image-level anomaly detection.
@@
3.3 Ablation Study
Our ablation study (Table 2) validates the individual and synergistic contributions of our two core components against a baseline using a linear adapter and static fusion. Integrating the Conv-LoRA Adapter alone boosts the pixel-level AUROC by 6.8%, confirming its critical role in providing the local inductive biases necessary for dense prediction. Separately, adding the Dynamic Fusion Gateway (DFG) improves the image-level AUROC by 4.7%, demonstrating the effectiveness of its vision-guided, adaptive fusion mechanism. The full ACD-CLIP model, which combines both, achieves the highest performance, confirming a powerful synergistic effect between enhancing feature representation and enabling dynamic cross-modal fusion.